# Neutron Star with Dark Matter Admixture: A Candidate for Bridging the Mass Gap

### v1

@@ Abstract
Neutron stars and black holes are the after death remnants of massive stars. However, according to the most recent observations, the neutron stars maximum mass is between $2.0-2.5\ M_{\odot}$ while black holes of less than 5 $M_{\odot}$ rarely has been observed. The region between the most massive neutron star and the least massive black hole is called the mass gap. If indeed its existence is confirmed by future observations, that indicates a gap in our understanding which seeks for explanation. In addition, the existence of compact objects within the mass gap should also be supported with the help of possible new theoretical scenarios. In this letter, we propose a possible explanation for the existence of compact objects within the mass gap region. Specifically, we propose that the mass gap region could be bridged by the existence of a hybrid compact object, composed of hadronic and self interacting - non annihilating fermionic dark matter, considering that the interaction between these two fluids it’s only gravitational. Fundamental questions about how these objects form and how they can be detected are also addressed.
Neutron Stars, Dark Matter, Mass Gap
Astrophysical compact objects come in two varieties: neutron stars (NSs) and black holes (BHs), which are the remnants of massive stars that reach the end of their lives with their cores collapsing in a supernova explosion [1, 2, 3]. NSs are physical nuclear laboratories consist mostly of neutrons, protons and electrons, whereas BHs are too heavy and dense described only by gravity. Both of these entities have uncertain mass limits. In a pioneering idea of Rhoades and Ruffini the optimum upper bound of mass of non-rotating NS was derived using a variations and conditions of standard General-Relativity equation of hydrostatic equilibrium, the Le Chattelier’s principle and the principle of causality [4]. This maximum mass of the equilibrium configuration of a NS cannot be larger than $3.2\ M_{\odot}$[4]. The problem of finding an optimum bound on the mass of non-rotating NS was considered also by other authors (see [5] and references therein) adopting similar assumptions to those of Ref. [4]. However, the majority of realistic nuclear matter equations of state equations predict that NSs maximum mass musdt be between $2.2-2.5\ M_{\odot}$ [1]. In addition, BHs of less than $5\ M_{\odot}$ rarely have been observed. These limits suggest a “mass gap” between the most massive NS and the least massive BH. The existence or not of the mass gap between NS and BH still remain an open problems. While some recent systematic studies support the existence of this gap, there are also recent observations of objects located within this area.
The first event were the mass gap issue started to get interesting was the GW190814 [6] reporting an object with mass $2.6\ M_{\odot}$, later another one, the GW190425 [7] reports a compact binary coalescence with total mass $3.4\ M_{\odot}$ and very recently, the GW230529 [8], came as an addition to the first observations estimating that the one of the components had a mass $3.6\ M_{\odot}$ supporting the possible existence of compact objects close to the lower limit of the mass gap. Also, a number of recent studies (see also Ref.[8]) that have found evidence for compact objects within the mass gap, include observations of non-interacting binary systems [9, 10], radio pulsar surveys [11], and Gravitational Waves (GWs) from compact binary coalescence’s [12, 13, 14, 15]. In any case, an identification whether the component in the gap is a NS or a BH would be remarkable. This distinction is crucial as it will help refine the boundaries of the mass gap, leading to positive advancements in our understanding of the physics of dense nuclear matter[16, 17, 18, 19, 20]. However, it has not yet been conclusively determined whether these components are black holes, neutron stars, or some other type of object. In both cases, one has to resolve this problem. In particular, if this gap really exists then a theoretical explanation for its existence must be given. On the other hand the possible existence of compact objects within this area must also be explained.
@@
Some compact objects observed in GW events have masses inside the gap between known NSs and BHs. The nature of these mass gap objects is unknown, as is the formation of their host binary systems (see also the relevant discussion in Ref.[8]). In fact, there are some theoretical suggestions, including the possibility of exotic compact objects, that could explain these components, such as gravastars [21], boson stars [22], or Planck-scale modifications of BH horizons [23, 24], which could also fall into this gap. Primordial black holes, which may have formed from over-dense regions in the early Universe, have also been proposed as potential candidates to fill the lower mass gap. [25]. In addition to astrophysical and primordial black holes within General Relativity (GR) and exotic compact objects, neutron stars predicted by alternative GR theories could also occupy this mass gap. For instance, in axionic scalar-tensor theory with viable phenomenological equations of state (EoS), such neutron stars could exist within this range. [26].
@@
In the present letter, we propose an alternative explanation involving compact objects within the mass gap region. More specifically, we consider that there could be a stable compact object consisting of NS coexistence with DM within the region in question. This idea, brings to mind the analogy of the DM in Galaxies as an admixture with the baryonic matter, playing a vital role to the rotation curve. Another important question is how DM could become mixed with ordinary matter in a NS. One well studied possibility is through capture as described in the [27, 28, 29, 30, 31, 32, 33, 34, 35]. Also many other possibilities of how celestial bodies like NSs can accumulate DM and interact with neutron matter - nucleons investigated in previous studies [36, 37, 38, 39]. Relevant discussion about the nature, the self-interaction of DM and dark objects are presented in Refs. [44, 41, 45, 46, 43, 47, 48, 42, 40, 50, 51, 49, 52, 53]
@@
In view of the above, a critical and logical question emerges: How can we distinguish these hybrid objects? What sets them apart from a classical NS or a pure DM star? Could the detection of associated gravitational waves (such as through tidal deformability) aid in their identification? The formation and therefore the existence of these objects, could be justified and only astrophysical observations remains to confirm or refute their existence. We also propose some possible ways that may lead to their observational identification.
We consider a compact hybrid object made by two fluids non-self annihilating DM particles admixed with NS matter (including mainly neutrons and a small fraction of protons and electrons). We consider also that the total matter, described by these two fluids, interact each other only gravitationally. In order to predict the bulk properties of the objects consisting of the aforementioned two fluids, one has to solve the coupled Tolman-Oppenheimer-Volkoff (TOV) equations, two for each fluid simultaneously. In the present work, in order to describe the NS matter, we use the EoS developed by Wang et al. [54] in the framework of the Relativistic Brueckner-Hartree-Fock (RBHF) theory. This is an EoS with microscopic origin and its predictions are in very good agreement with both the measured maximum masses and some astrophysical constraints for radii [55]. Now, concerning the DM particles, we consider that they are relativistic fermions, which interact each other through a repulsive force. We consider Yukawa type interaction [40]:
@@
where ${\rm g}_{\chi}$ is the coupling constant and $m_{\phi}$ is the mediator mass. The role of interaction is decisive for the creation and stability of the mass gap objects. As we mentioned before, the DM is a self-interacting Fermi gas, where the contribution on the energy density of the self-interaction is given by $\frac{y^{2}}{2}(\hbar c)^{3}n_{\chi}^{2}$, where $n_{\chi}$ is the DM density and $y={\rm g}_{\chi}/m_{\phi}c^{2}$ (in units MeV-1), is the interaction strength. The interaction strength $y$ can be constrained by observational limits. These limits imposed on the cross section of the self-interaction [56, 57]. In particular according to [58, 59, 60, 61] it holds $\sigma/m_{\chi}\sim 0.1-10\ {\rm cm}^{2}/{\rm g}$. Moreover, it has been showed that the Born approximation is very accurate in the region $m_{\chi}\sim 1\ {\rm GeV}~{}$[43, 42, 62]. Thus, we found that the interaction parameter $y$ approximately differs across a range spanning two orders of magnitude $\sim(0.001-0.1)({\rm GeV}/m_{\chi}c^{2})^{1/4}({\rm MeV}^{-1})$ [56].

### v2

@@ Abstract
Neutron stars, white dwarfs and black holes are the after death remnants of massive stars. However, according to the most recent observations, the neutron stars maximum mass is between $2.0-2.5\ M_{\odot}$ while black holes of less than 5 $M_{\odot}$ has not yet been observed. The region between the most massive neutron star and the least massive black hole is called the mass-gap. If indeed its existence is confirmed by future observations, that indicates a gap in our understanding which seeks for explanation. In addition, the existence of compact objects within the mass-gap should also be supported with the help of possible new theoretical scenarios. In this study, we propose a possible explanation for the existence of compact objects within the mass-gap region. Specifically, we propose that the mass-gap region could be bridged by the existence of a hybrid compact object, composed of hadronic and self interacting and non-annihilating fermionic dark matter, considering that the interaction between these two fluids it’s only gravitational. Fundamental questions about how these objects form and how they can be detected are also addressed.
keywords:
Neutron stars; dark matter; mass gap.
[…]
(Day Month Year)
1 Introduction
Astrophysical compact objects come in three varieties: neutron stars (NSs), white dwarfs (WDs) and black holes (BHs), which are the remnants of massive stars that reach the end of their lives with their cores collapsing in a supernova explosion [1, 2, 3]. NSs are physical nuclear laboratories with complex internal composition with main components neutrons, protons and electrons but also some exotic particles including hyperons, pions and kaons, muons or even deconfined quark matter. BHs, which are formed when massive stars collapse under their own gravity, are objects with a very strong gravity characterized mainly by the three quantities that is the mass, the spin and the charge. Both of these entities, NSs and BHs, have uncertain mass limits.
In a pioneering idea of Rhoades and Ruffini an upper bound of mass of non-rotating NS was derived using a variations and conditions of standard General-Relativity equation of hydrostatic equilibrium, the Le Chattelier’s principle and the principle of causality [4]. This maximum mass of the equilibrium configuration of a NS cannot be larger than $3.2\ M_{\odot}$[4]. The problem of placing an upper bound on the mass of non-rotating NS was considered also by other authors (see [5] and references therein) adopting similar assumptions to those of Ref. [4]. However, the majority of nuclear matter equations of state (EoS), and which have been developed based on all possible assumptions regarding the composition of NSs, predict that their maximum mass must be between $2.2-2.5\ M_{\odot}$ [1]. Moreover, the authors in Ref. [6], combining the gravitational waves (GW) observations of merging systems of binary neutron stars (exploiting the recent observation of the GW event GW170817 [7] and drawing from basic arguments on kilonova modeling of GRB 170817A) and quasi-universal relations, set constraints on the maximum mass that can be attained by nonrotating stellar models of neutron stars. More specifically, they set limits for the maximum mass to be $2.01_{-0.04}^{+0.04}\leq M_{\rm max}/M_{\odot}\lesssim 2.16_{-0.15}^{+0.17}$. In Ref. [8] the authors revisited the constraints on the maximum mass of cold spherical neutron stars coming from the observational results of GW170817 and they found that the maximum mass of cold spherical neutron stars can be only weakly constrained as $M_{\rm max}\lesssim 2.3\ M_{\odot}$. In the same framework, the authors in Ref. [9] used the GW170817 event by adopting the scenario in which a hypermassive compact star remnant formed in a merger evolves into a supramassive compact star that collapses into a black hole once the stability line for such stars is crossed. They deduced an upper limit on the maximum mass of static, cold neutron stars $2.15_{-0.17}^{+0.18}\leq M_{\rm max}/M_{\odot}\leq 2.24_{-0.44}^{+0.45}$ for the typical range of entropy per baryon, $2\leq S/A\leq 3$ and electron fraction $Y_{e}=0.1$ characterizing the hot hypermassive star.
In addition, BHs of less than $5\ M_{\odot}$ has not yet been observed. These limits suggest a “mass gap” between the most massive NS and the least massive BH. More specifically, in the study [10], the authors, having conducted an extensive study, concluded that there is a paucity of BHs with masses $\simeq 2-5\ M_{\odot}$. Their results confirmed and strengthen an earlier finding of Bailyn et al. [11], who argued for a low-mass gap based on a limited sample of BHs. Moreover, Farr et al. [12] examine the existence of a gap between the most massive NSs and the least massive BHs by considering the value, $M_{1\%}$, of the $1\%$ quantile from each BH mass distribution as the lower bound of BH masses. Their analysis generates posterior distributions for $M_{1\%}$ where the best model (the power law) fitted to the low-mass systems has a distribution of lower bounds with $M_{1\%}>4.3\ M_{\odot}$ with $90\%$ confidence, while the best model (the exponential) fitted to all 20 systems has $M_{1\%}>4.5\ M_{\odot}$ with $90\%$ confidence. They concluded that their sample of BH masses offers compelling evidence of a gap between the maximum NS mass and the lower limit of BH masses [12]. On the other hand however, in Ref. [13] the authors model the dense globular cluster NGC 1851 with a grid of similar dense star clusters utilizing the state-of-the-art Monte Carlo N-body code Cluster Monte Carlo, and they specifically studied the formation of LMG-BHs. They demonstrated that both massive star evolution and dynamical interactions can play a role in the formation of low-mass black holes. It is worth to mention here that in the analysis of the multimessenger event GW170817–GRB170817A, in Ref. [14] it has been identified a BH of $2.8\ M_{\odot}$. If confirmed, the proposed mass gap will be narrowed to $2.4-2.8\ M_{\odot}$, assuming $2.4\ M_{\odot}$, as the upper limit for the mass of a NS. Furthermore, it is believed that the mass gap mentioned above could become more accurate with future observations of double neutron star (DNS) mergers, due to the variability in the mass distribution of DNS systems. [15].
In any case, the existence or not of the mass gap between NSs and BHs still remain an open problems. While some recent systematic studies support the existence of this gap, there are also recent observations of objects located within this area. The first event were the mass gap issue started to get interesting was the GW190814 [16] reporting an object with mass $2.6\ M_{\odot}$, later another one, the GW190425 [17] reports a compact binary coalescence with total mass $3.4\ M_{\odot}$. Very recently, the authors in Ref. [18] (GW230529 event), report the observation of a coalescing compact binary with component masses $2.5–4.5\ M_{\odot}$ and $1.2–2.0\ M_{\odot}$ (all measurements quoted at the $90\%$ credible level). The discovery of this system suggests an increase in the anticipated rate of NS–BH mergers accompanied by electromagnetic counterparts and adds further support to the existence of compact objects within the supposed lower mass gap. Also, a number of recent studies (see also Ref. [19]) that have found evidence for compact objects within the mass gap, include observations of non-interacting binary systems [20, 21], radio pulsar surveys [22], and Gravitational Waves (GWs) from compact binary coalescence’s [16, 23, 7, 24].
Regardless of the situation, an identification whether the component in the gap is a NS or a BH would be remarkable. This distinction is crucial as it will help refine the boundaries of the mass gap, leading to positive advancements in our understanding of the physics of dense nuclear matter [6, 11, 12, 10, 25, 26]. However, it has not yet been conclusively determined whether these components are BHs, NSs, or some other type of object. Future, more detailed observations may potentially shed light on which of these objects.
@@ 1 Introduction
In particular, if this gap truly exists, a theoretical explanation for its presence must be provided. On the other hand, the potential existence of compact objects within this region must also be accounted for. Some compact objects observed in gravitational wave events have masses within the gap between known NSs and BHs. The nature of these mass-gap objects is still unknown, as is the formation process of their host binary systems (see also the related discussion in Ref. [19]). In fact, there are some theoretical suggestions, including the possibility of exotic compact objects, that could explain these components, such as gravastars [27], boson stars [28], or Planck-scale modifications of BH horizons [29, 30], which could also fall into this gap.
Primordial BHs, which may have formed from over-dense regions in the early Universe, have also been suggested as potential candidates to populate the lower mass gap [31]. In addition to astrophysical and primordial BHs within General Relativity (GR) and exotic compact objects, NSs predicted by alternative theories of GR could also occupy this mass gap. For example, in axionic scalar-tensor theory with viable phenomenological equations of state (EoS), such NSs could fall within this range [32].
@@
In the present work, we propose an alternative explanation involving compact objects within the mass gap region. More specifically, we consider that there could be a stable compact object consisting of NS coexistence with DM within the region in question. This idea, brings to mind the analogy of the DM in Galaxies as an admixture with the baryonic matter, playing a vital role to the rotation curve. Another significant question is how DM could integrate with ordinary matter in a NS. A well-explored possibility is through capture, as detailed in Refs. [33, 34, 35, 36, 37, 38, 39, 40, 41]. Numerous other possibilities for how celestial bodies, such as NSs, can accumulate DM and interact with neutron matter—nucleons have been explored in previous studies [42, 43, 44].
In particular in Ref. [42] the authors consider that NSs near the Galactic center may exist in a dense DM environment and could accumulate local DM clouds if it interacts efficiently with neutrons. This makes them valuable for studying DM annihilation. In models with light mediators, DM may annihilate into mediators that escape the star and subsequently decay into neutrinos outside of it. The research assesses constraints on heavy DM theories (TeV to PeV) involving long-lived mediators that decay into neutrinos, using high-energy neutrino observatory data and future projections These observations complement terrestrial searches and explore otherwise unreachable DM parameter spaces [42].
In Ref. [43], the authors suggest that DM from the Galactic halo can accumulate in NSs and potentially convert them into BHs with masses under $2.5\ M_{\odot}$, assuming the DM particles are heavy, stable, and interact with nucleons. They show that the lack of gravitational wave detections from mergers involving such low-mass BHs can be used to place constraints on the interactions between non-annihilating DM particles and nucleons.
The authors in Ref. [44] explore a model of composite DM involving stable particles with a charge of 2-2 that bind with primordial helium ($He$) nuclei through Coulomb forces to form $OHe$ atoms. They investigate how these dark atoms could be captured by regular matter, suggesting the potential existence of stable, superheavy nuclei enriched with $O$ and $O$ nuclearites, where heavy $O$ DM fermions are electromagnetically bound with ordinary nuclear matter. The study also examines how the accumulation of $OHe$ atoms in stars might influence stellar evolution, broadening the range of indirect methods to probe composite DM [44]. Relevant discussion about the nature, the self-interaction of DM and dark objects are presented also in Refs. [49, 46, 50, 51, 48, 52, 53, 47, 45, 55, 56, 54, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70].
@@
In view of the above, a critical and logical question emerges: How can we distinguish these hybrid objects? What sets them apart from a classical NS or a pure DM star? Could the detection of associated gravitational waves (such as through tidal deformability) aid in their identification? The formation and therefore the existence of these objects, could be justified and only astrophysical observations remains to confirm or refute their existence. In the present study, we also propose some possible ways that may lead to their observational identification.
The paper is organized as follows: In Section II, we introduce the theoretical model which encompasses the equations of state for both DM and NS matter, as well as the stability conditions for the hybrid objects. In Section III we present and discuss the results of this study while in Section V, we conclude with our final remarks.
2 The theoretical model
2.1 Two fluid model
We consider a compact hybrid object made up of two fluids: non-self-annihilating DM particles mixed with NS matter. The composition of NSs is assumed to be in the hadronic phase, predominantly consisting of neutrons, with a small proportion of protons and electrons. Additionally, we assume that the two fluids, which together describe the total matter, interact exclusively via gravity. To predict the macroscopic properties of objects composed of these two fluids, it is necessary to solve the coupled Tolman-Oppenheimer-Volkoff (TOV) equations, with two equations for each fluid, simultaneously. These four equations are outlined below [72, 71]
$\frac{dP_{\rm NS}(r)}{dr}$ $=$ $-\frac{G{\cal E}_{\rm NS}(r)M(r)}{c^{2}r^{2}}\left(1+\frac{P_{\rm NS}(r)}{{\cal E}_{\rm NS}(r)}\right)$ (1)
$\times$ $\left(1+\frac{4\pi P(r)r^{3}}{M(r)c^{2}}\right)\left(1-\frac{2GM(r)}{c^{2}r}\right)^{-1}$
$\frac{dM_{\rm NS}(r)}{dr}=\frac{4\pi r^{2}}{c^{2}}{\cal E}_{\rm NS}(r)$ (2)
$\frac{dP_{\rm DM}(r)}{dr}$ $=$ $-\frac{G{\cal E}_{\rm DM}(r)M(r)}{c^{2}r^{2}}\left(1+\frac{P_{\rm DM}(r)}{{\cal E}_{\rm DM}(r)}\right)$ (3)
$\times$ $\left(1+\frac{4\pi P(r)r^{3}}{M(r)c^{2}}\right)\left(1-\frac{2GM(r)}{c^{2}r}\right)^{-1}$
$\frac{dM_{\rm DM}(r)}{dr}=\frac{4\pi r^{2}}{c^{2}}{\cal E}_{\rm DM}(r)$ (4)
where the subscripts $\rm NS$ and $\rm DM$ stand for the NS and DM respectively. Moreover $P_{i}(r)$, ${\cal E}_{i}(r)$, $M_{i}(r)$ ($i={\rm NS},{\rm DM}$) are the specific values of pressure, energy density and mass. Finally, the total pressure, energy density and mass are given by $P(r)=P_{\rm NS}(r)+P_{\rm DM}(r)$, ${\cal E}(r)={\cal E}_{\rm NS}(r)+{\cal E}_{\rm DM}(r)$ and $M(r)=M_{\rm NS}(r)+M_{\rm DM}(r)$ respectively. To solve the TOV equations mentioned above, it is essential to know the equations of state that characterize both NS matter (hadronic matter) and DM. The next two subsections focus on this topic.
2.2 Hadronic Equation of State
In this study, to describe NS matter, we employ the equation of state developed by Wang et al. [73] within the framework of Relativistic Brueckner-Hartree-Fock (RBHF) theory. This EoS has a microscopic foundation and its predictions align well with both the observed maximum masses and certain astrophysical constraints on radii [73, 74].
To be more specific, the NS Interior Composition Explorer (NICER) mission recently reported two independent Bayesian parameter estimates for the mass and equatorial radius of the millisecond pulsar PSR J0030+0451 [75]. The predictions from the RBHF theory are fully consistent with the constraints provided by NICER [73]. Furthermore, the maximum NS masses $M_{\rm max}$ obtained from the RBHF theory align with the available astrophysical constraints from observations of massive NSs, such as PSR J1614-2230 [76, 77], PSR J0348+0432 [78], and PSR J0740+6620 [79, 80].
2.3 Dark matter model
Regarding the DM particles, we assume they are relativistic fermions that interact with each other through a repulsive force. We consider a Yukawa-type interaction for this purpose [45]
@@ 2.3 Dark matter model
where ${\rm g}_{\chi}$ and $m_{\phi}$ are the coupling constant and the mediator mass respectively. The interaction plays a crucial role in the formation and stability of mass gap objects. The contribution on the energy density of the self-interaction is given by
${\cal E}_{SI}(n_{\chi})=\frac{y^{2}}{2}(\hbar c)^{3}n_{\chi}^{2}$ (6)
where $y={\rm g}_{\chi}/m_{\phi}c^{2}$ (in units MeV-1). In this case the total energy density of the DM particles is given by [55]
${\cal E}_{\rm DM}(n_{\chi})$ $=$ $\frac{(m_{\chi}c^{2})^{4}}{(\hbar c)^{3}8\pi^{2}}\left[x\sqrt{1+x^{2}}(1+2x^{2})\right$. (7)
- $\left.\ln(x+\sqrt{1+x^{2}})\right]+\frac{y^{2}}{2}(\hbar c)^{3}n_{\chi}^{2}$
The corresponding pressure is calculated straightforward by using the definition
$P_{\rm DM}(n_{\chi})=n_{\chi}\frac{d{\cal E}_{\rm DM}(n_{\chi})}{dn_{\chi}}-{\cal E}_{\rm DM}(n_{\chi})$ (8)
and is given by
$P_{\rm DM}(n_{\chi})$ $=$ $\frac{(m_{\chi}c^{2})^{4}}{(\hbar c)^{3}8\pi^{2}}\left[x\sqrt{1+x^{2}}(2x^{2}/3-1)\right$. (9)
$+$ $\left.\ln(x+\sqrt{1+x^{2}})\right]+\frac{y^{2}}{2}(\hbar c)^{3}n_{\chi}^{2}$
where $m_{\chi}$ is the particle mass and
$x=\frac{(\hbar c)(3\pi^{2}n_{\chi})^{1/3}}{m_{\chi}c^{2}}.$
The total energy density and pressure of the two-fluid mixing, considering that the contribution of NS matter are defined as ${\cal E}_{\rm NS}(n_{b})$ and $P_{\rm NS}(n_{b})$ where both are functions of the baryon density $n_{b}$, are given simply by the following sums
${\cal E}(n_{b},n_{\chi})={\cal E}_{\rm NS}(n_{b})+{\cal E}_{\rm DM}(n_{\chi})$ (10)
$P(n_{b},n_{\chi})=P_{\rm NS}(n_{b})+P_{\rm DM}(n_{\chi}).$ (11)
@@
As we mentioned before, the DM is a self-interacting Fermi gas, where the contribution on the energy density of the self-interaction is given by $\frac{y^{2}}{2}(\hbar c)^{3}n_{\chi}^{2}$, where $n_{\chi}$ is the DM density and $y={\rm g}_{\chi}/m_{\phi}c^{2}$ (in units MeV-1), is the interaction strength. The interaction strength $y$ can be constrained by observational limits. These limits imposed on the cross section of the self-interaction [81, 82]. In particular, according to [83, 84, 85, 86] it holds $\sigma/m_{\chi}\sim 0.1-10\ {\rm cm}^{2}/{\rm g}$. Moreover, it has been showed that the Born approximation is very accurate in the region $m_{\chi}\sim 1\ {\rm GeV}$~[48, 47, 87]. Thus, we found that the interaction parameter $y$ varies in the range $\sim(0.001-0.1)({\rm GeV}/m_{\chi}c^{2})^{1/4}({\rm MeV}^{-1})$ [81].
2.4 Tidal deformability
A crucial source for gravitational wave detectors comes from the gravitational waves generated during the late phase of the inspiral of a binary NS system, just before the merger [88, 89, 90]. This kind of source leads to the measurement of various properties of NSs. In the inspiral phase the tidal effects can be detected [89].
The $k_{2}$ parameter, also known as tidal Love number, depends on the equation of state and describes the response of a NS to the tidal field $E_{ij}$ [89]. The exact relation is given below
$Q_{ij}=-\frac{2}{3}k_{2}\frac{R^{5}}{G}E_{ij}\equiv-\lambda E_{ij},$ (12)
where $R$ is the radius of the compact object and $\lambda=2R^{5}k_{2}/3G$ is the tidal deformability. The tidal Love number $k_{2}$ is given by [89, 90]
$k_{2}$ $=$ $\frac{8\beta^{5}}{5}\left(1-2\beta\right)^{2}\left[2-y_{R}+(y_{R}-1)2\beta\right]$ (13)
$\times$ $\left[\frac{}{}2\beta\left(6-3y_{R}+3\beta(5y_{R}-8)\right)\right$.
$+$ $4\beta^{3}\left.\left(13-11y_{R}+\beta(3y_{R}-2)+2\beta^{2}(1+y_{R})\right)\frac{}{}\right$.
$+$ $\left.3\left(1-2\beta\right)^{2}\left[2-y_{R}+2\beta(y_{R}-1)\right]{\rm ln}\left(1-2\beta\right)\right]^{-1},$
where $\beta=GM/Rc^{2}$ is the compactness of the star. The parameter $y_{R}$ is determined by solving numerically the following differential equation (for a detailed analysis for the case of the two fluid model see also Ref. [91])
$r\frac{dy(r)}{dr}+y^{2}(r)+y(r)F(r)+r^{2}Q(r)=0.$ (14)
$F(r)$ and $Q(r)$ are functionals of the total energy density ${\cal E}(r)$, total pressure $P(r)$, and total mass $M(r)$ defined as [88, 91]
$F(r)=\left[1-\frac{4\pi r^{2}G}{c^{4}}\left({\cal E}(r)-P(r)\right)\right]\left(1-\frac{2M(r)G}{rc^{2}}\right)^{-1},$ (15)
and
$r^{2}Q(r)$ $=$ $\frac{4\pi r^{2}G}{c^{4}}\left[\frac{}{}5{\cal E}(r)+9P(r)\right$. (16)
$+$ $\left.\frac{{\cal E}_{\rm NS}(r)+P_{\rm NS}(r)}{\partial P_{\rm NS}(r)/\partial{\cal E}_{\rm NS}(r)}+\frac{{\cal E}_{\rm DM}(r)+P_{\rm DM}(r)}{\partial P_{\rm DM}(r)/\partial{\cal E}_{\rm DM}(r)}\right]$
$\times$ $\left(1-\frac{2M(r)G}{rc^{2}}\right)^{-1}-6\left(1-\frac{2M(r)G}{rc^{2}}\right)^{-1}$
- $\frac{4M^{2}(r)G^{2}}{r^{2}c^{4}}\left(1+\frac{4\pi r^{3}P(r)}{M(r)c^{2}}\right)^{2}\left(1-\frac{2M(r)G}{rc^{2}}\right)^{-2}.$
Eq. (14) must be solved numerically and self-consistently with the coupled TOV equations, which correspond to the two-fluid model, under the following conditions $y(0)=2$, $M_{NS}(0)=0$ and $M_{DM}(0)=0$ and with initial values of pressure the pair $\{P_{c}^{\rm NS},P_{c}^{\rm DM}$[88, 90, 91]. From the numerical solution of TOV equations, the neutron and DM masses and radii can be computed, while the corresponding solution of the differential Eq. (14) provides the value of $y_{R}=y(R)$. The last parameter along with the quantity $\beta$ are the basic ingredients of the tidal Love number $k_{2}$. Additionally, a related quantity in our study is the dimensionless tidal deformability $\Lambda$ defined as
$\Lambda=\frac{2}{3}k_{2}\left(\frac{c^{2}R}{GM}\right)^{5}.$ (17)
$\Lambda$ is a key parameter that can be determined through the detection of the associated gravitational waves and can provide valuable insights into both the structure and size of compact stars [88, 90, 91].
2.5 Stability Conditions
Ultimately, the essential research required involves examining the stability of these hybrid objects. Only after confirming their stability can we hypothesize that they may exist in nature. A common method for assessing stability is to analyze small radial perturbations of the equilibrium state by solving the Sturm-Liouville eigenvalue equation, which yields the eigenfrequencies $\omega_{n}$ [1]. These eigenfrequencies form a discrete series where $\omega_{n}^{2}<\omega_{n+1}^{2}$ $n=0,1,2\cdots$ where $\omega_{n}$ being real numbers [92]. If $\omega_{n}^{2}$ is negative, the radial perturbation grows exponentially, leading to the collapse of the star. The star is considered stable only when all eigenfrequencies are positive [93] (see also the detailed discussion in Ref. [1]). A notable analysis of pulsation equations for compact stars consisting of multiple perfect fluids was recently conducted in Ref. [94, 91].